5. Solne 2 dim we, Lo-A) uzulx,y,t) 2 + 3 2 1/2 = 24 460, y, t) =0 ulx, o, t=0 ultry, t) =0 u(x, It)=0 Lux,y, O) = xy+x,
Answer #1

Answer U = Uxx + Ugy. u coy, t)=0, U14,4, t) =0., 4601,0,1)=0, ube,8,4)0 u (x, y,0)= xy + y² , ut (x, y,00=9(x,y) Assume u(3,2 Console - SILV) 0 => x 10 =0 - Slx,0) =0 = Y(0)=0 --> 5 CX,k)=0 ->y (k)=0 consider x+ MX =0, x(0)-0, X(0)=0 xn () = sin (PAllxy,t) = sin(TX) sin(many) cos (fmt)+Bom sin (samo)) u(1,9,0) = xy + y2 op z 3 sin () si (TY) (Am D-xygge U DE => , Aponsod